OpenStudy (anonymous):

given f(x) = (2x-2)/4 solver for f^1(3)

OpenStudy (anonymous):

:/

OpenStudy (anonymous):

@amistre64 @phi @jim_thompson5910 @cwrw238

OpenStudy (amistre64):

when y=3, what does x have to be? or, for what value of x, does f(x) = 3

OpenStudy (amistre64):

\[f(x)=3\] \[f^{-1}(f(x))=f^{-1}(3)\] \[x=f^{-1}(3)\]

OpenStudy (anonymous):

I know that.... Solving it is my issue because I can't just plug it in

OpenStudy (amistre64):

sure you can, i just demonstarted that f(x) = 3

OpenStudy (amistre64):

f(x) = (2x-2)/4 3 = (2x-2)/4

OpenStudy (amistre64):

the value of x is equal to f^-1(3)

OpenStudy (amistre64):

even if we dont go that route, but replace x by f^-1(3) \[f(x)=\frac{2x-2}{4}\] \[f(f^{-1}(3))=\frac{2(f^{-1}(3))-2}{4}\] \[3=\frac{2(f^{-1}(3))-2}{4}\] solve for f^(-1)(3)

OpenStudy (anonymous):

If I plug it in, I get 1. i don't know what the ^-1 means

OpenStudy (amistre64):

^-1 is inverse notation ....

OpenStudy (anonymous):

I know that

OpenStudy (amistre64):

youre giving me mixed results here ... you either know or you dont; it cant be both :/

OpenStudy (anonymous):

Let me rephrase. I KNOW what the ^-1 IS but I don't know what to do with it

OpenStudy (amistre64):

hmm, you recall that y = f(x) right?

OpenStudy (anonymous):

yep, sorry for the miscommunication

OpenStudy (amistre64):

to undo the f, we apply its inverse; so f^-1 each side \[y=f(x)\] \[f^{-1}(y)=f^{-1}(f(x))\] \[f^{-1}(y)=x\] letting y = 3 \[f^{-1}(3)=x\] and since y = f(x), let f(x) = 3 and solve for x

OpenStudy (amistre64):

hmm, you recall that y = f(x) right?

OpenStudy (anonymous):

so the Y actually equals 3? Although I don't have a Y? Seems weird that they'd give me that

OpenStudy (amistre64):

f(x) is a curve in the xy plane; for some value of x, we can measure the value of the curve, and relate it to the y axis.

OpenStudy (phi):

you could think like this f(x) is a rule that changes x into y the inverse goes the other way, given y, you find x

OpenStudy (amistre64):